# Make Every Penny Count: Difficulty-Adaptive Self-Consistency for Cost-Efficient Reasoning

### v2

@@ Abstract
Self-consistency (SC), a widely used decoding strategy for chain-of-thought reasoning, shows significant gains across various multi-step reasoning tasks but comes with a high cost due to multiple sampling with the preset size. Its variants, Adaptive self-consistency (ASC) and Early-stopping self-consistency (ESC), dynamically adjust the number of samples based on the posterior distribution of a set of pre-samples, reducing the cost of SC with minimal impact on performance. Both methods, however, do not exploit the prior information about question difficulty. It often results in unnecessary repeated sampling for easy questions that could be accurately answered with just one attempt, wasting resources. To tackle this problem, we propose Difficulty-Adaptive Self-Consistency (DSC), which leverages the difficulty information from both prior and posterior perspectives to adaptively allocate inference resources, further reducing the cost of SC. To demonstrate the effectiveness of DSC, we conduct extensive experiments on three popular categories of reasoning tasks: arithmetic, commonsense and symbolic reasoning on six benchmarks. The empirical results show that DSC consistently surpasses the strong baseline ASC and ESC in terms of costs by a significant margin, while attaining comparable performances.
@@
Xinglin Wang1, Shaoxiong Feng2, Yiwei Li1, Peiwen Yuan1, Yueqi Zhang1, Boyuan Pan2, Heda Wang2, Yao Hu2, Kan Li1222Corresponding author. 1 School of Computer Science, Beijing Institute of Technology 2 Xiaohongshu Inc {wangxinglin,liyiwei,peiwenyuan, zhangyq, likan}@bit.edu.cn {shaoxiongfeng2023,whd.thu}@gmail.com {panboyuan,xiahou}@xiaohongshu.com
@@ 1 Introduction
However, both approaches still suffer from the following two shortcomings: (1) Both require a certain amount of pre-sampling for all problems, which still results in redundant waste. As shown in Figure 1, there is a high overlap of inference results between SC and CoT, which suggests that only a small portion of questions (3.8% for GPT-4) benefit from SC. Generating multiple samples for the remaining questions compared to CoT (only sampling once) results in a significant waste of costs. Although ASC and ESC somewhat reduce the ineffective cost of SC by decreasing average sample size, there remains redundant sampling for simple problems111At least 3 samples per problem for ASC and 4 for ESC.. (2) Multiple re-samples bring additional significant input costs. Both ASC and ESC focus solely on reducing the number of outputs, without considering the extra input costs brought by multiple re-samples (Table 1). Consequently, for problems requiring multiple sampling times, ASC and ESC will introduce substantial extra input costs, sometimes outweighing the savings achieved through output sampling reduction (Table 3).
To alleviate these issues, we take inspiration from the strategies humans employ when solving reasoning problems within a limited total time. As illustrated in Figure 2, humans pre-assess the difficulty of the problems before reasoning, and adaptively allocating problem-solving time based on the accessed difficulty. In light of this, we propose Difficulty-Adaptive Self-Consistency (DSC), a novel method which leverages the difficulty information from both prior and posterior perspectives to further reduce the inference computational cost of SC. As illustrated in Figure 3, DSC consists of three steps. Firstly, we propose Difficulty Ranking algorithm which utilizes the LLM itself to rank the difficulty of the given problem set222Although using LLM for difficulty ranking will introduce additional cost, our later experiments will prove that this part of the cost is totally acceptable.. Based on this, we propose a Problem Partition strategy that divides the problem set into two parts, easy and hard, using the information regarding difficulty rankings. For problems belonging to easy part, a single CoT sampling is performed, which saves cost without sacrificing performance. Lastly, we propose Sample Size Pre-Allocation algorithm to predict the sample size needed for problems belonging to hard part, which reduce the number of re-samples and save the cost of inputs.
@@
We conducted extensive experiments on six popular benchmarks and validated the effectiveness of our proposed method.
@@ 2.1 Difficulty Ranking
As problems usually do not carry difficulty label themselves, we seek to utilize the powerful comparative and ranking capabilities of LLM to obtain the difficulty rank of the given problem set. Considering the limited context window size of LLM and its lost in the middle (Liu et al., 2024) issue in understanding long contexts, we randomly divide the problem set of into batches of size () and let LLM rank the difficulty of the problems in each batch333See Appendix A.1 for corresponding prompts.. Since a single random split only allows us to obtain the difficulty ranking of each problem within the corresponding batch, we perform times random batch splitting, and sort the average difficulty rankings of each problem in different batches to obtain the difficulty ranking of entire problem set. Algorithm 1 illustrates the specific implementation of difficulty ranking.
@@ 2.2 Problem Partition
After obtaining the problem set with difficulty ranking, we aim to find which part of the problems only require LLM to perform CoT sampling once. A simple and intuitive idea is that when LLM is very confident about a number of continuous problems sorted by difficulty (the results of all pre-samples are the same), it can be inferred that the problems easier than these problems only need one CoT sampling. Guided by this, we design the Problem Partition algorithm, illustrated in Algorithm 2. Specifically, we pre-sample the problem set sorted by difficulty from hard to easy, and store the entropy of pre-sampling results of each problem in a list. We stop pre-sampling when the latest items in the list are all zero. The remaining problems without pre-sampling are then divided into the easy part, and a single CoT sampling is performed for each.
@@ 2.3 Sample Size Pre-Allocation
After performing CoT sampling on the easy part problems, we seek to predict and allocate the sample size for the problems belonging to hard part, so as to mitigate the substantial cost brought by multiple re-samples. Considering that the required sample size for each problem should be similar to those needed for problems of comparable difficulty, we predict the sample size of the current problem based on the total sample size of the nearest easier444As the pre-allocated sample size could be larger than the required sample size (causing higher output costs), we use the sample size of easier rather than harder neighboring problems for prediction. problems. Algorithm 3 shows the workflow of Sample Size Pre-Allocation. Specifically, we sample questions belonging to the hard part from easy to difficult. For the current question to be inferred, we predict its pre-allocation sample size based on the average total sample size of its previous questions. Then, we judge the distribution of the current samples based on the stopping criteria 555Following ASC, we use Dirichlet Stopping Criteria. Please refer to Appendix A.6 for more details.. When the criteria is not met, we re-sample based on the expansion window , until the sampling distribution meets the criteria for stopping sampling or the number of samples reaches the max sample size . After the sampling for the current question ends, we add its samples and total sample size to the and lists respectively, in order to pre-allocate the sample size for the next question.
@@ 3.2 Main Results
Table 2 summarizes the cost and accuracy of SC, ASC, ESC, and proposed DSC for each dataset. We show that DSC consistently outperforms all baselines on cost by a significant margin across all datasets, while barely affecting performance. Specifically, DSC reduces the cost on GPT-4 by an average of 65.29% and on GPT-3.5-Turbo by 56.04% compared to SC. In comparison to the strongest baseline method ESC, DSC reduces the cost on GPT-4 by an average of 24.81% and on GPT-3.5-Turbo by 21.86%, which demonstrates the effectiveness of DSC.
@@
| Model | Metric | Unmixed | Mixed |
| GPT-4 | Spearman | 0.6451 | 0.5148 |
| Pearson | 0.6487 | 0.5197 | |
| Kendall | 0.5060 | 0.3931 | |
| GPT-3.5-Turbo | Spearman | 0.4828 | 0.4070 |
| Pearson | 0.4887 | 0.4145 | |
| Kendall | 0.3697 | 0.3073 |
@@ 3.3.1 Difficulty Ranking
As we instruct LLM to produce extremely concise outputs for Difficulty Ranking, the cost of output tokens brought by Difficulty ranking is very low777With an average consumption of 15 tokens for each question.. Table 6 gives comparison of input tokens brought by Difficulty Ranking and few-shot reasoning across six datasets. We show that the average input cost for each problem’s difficulty ranking is less than the average single input cost for its few-shot reasoning. Considering that ESC and ASC require multiple re-samples for reasoning (multiple few-shot inputs), this additional input cost of Difficulty Ranking is totally acceptable.
Mixing types of questions lowers the effectiveness of Difficulty Ranking.
To explore whether Difficulty Ranking can effectively rank problems of different types in terms of difficulty, we compare the performance of Difficulty Ranking under mixed and unmixed problem types on MATH, as shown in Table 7. Specifically, for unmixed, we perform Difficulty Ranking on seven subsets separately; for mixed, we directly execute Difficulty Ranking on random shuffled MATH dataset, and calculate the correlation on the entire dataset for both. The results indicate that mixing different types of questions for LLM to rank can lead to a decrease in performance. This might be due to the challenge for LLM in adhering to the same assessment standards for various types of questions.888It could also lead to inaccurate ranking for humans.
@@ 3.3.2 Problem Partition
Regarding the intuitive idea proposed in Section 2.2, a straightforward question is how large the judge window size needs to be to ensure the accuracy on simpler problems with single CoT sampling. To answer this question, we conduct an experiment of DSC under different judge window size with GPT-4 on GSM8K. As shown in Figure 4, the accuracy increases with the rise of until it reaches 32. Meanwhile, the cost of DSC also increases with the enlargement of , as the easy part divided by the problem partition decreases with its increase. Therefore, it is a good choice to set a smaller judge window size while ensuring accuracy.999We select 32 as the default value for .
@@ 3.3.3 Sample Size Pre-Allocation
Given that Sample Size Pre-Allocation can trade output tokens for savings in input tokens, a natural question is when the max total savings can be achieved. As shown in Figure 5, we count the tokens of input and output and the corresponding cost of DSC under different prediction window size . The experimental results show that the cost of DSC decreases rapidly first and then slighly increases with the increase of , and the input tokens also follow a similar trend. This is because as increases, the sample size of the current question is predicted by more simpler questions. When the value of is large enough, the predicted sample size will gradually be less than its actual value, leading to the need for more re-samples. In contrast, the tokens of the output part gradually decrease, which is due to as m increases, the predicted sample size will continue to decrease, leading to a gradual decrease in the situation where the predicted sample size is greater than its actual need, saving the output tokens from redundant sampling. In summary, it is suggested to select within a relatively moderate range (e.g. from 8 to 32).
@@ 4 Related Work
Chain-of-thought prompting has been proven to be an effective method of solving complex reasoning problems (Wei et al., 2022a). By following the pattern of gradually solving sub-problems, few-shot CoT (Fu et al., 2023) are capable of stimulating LLM reasoning abilities. On this basis, Least-to-most prompting (Zhou et al., 2023) suggests explicitly splitting the problem and solving them step by step. Zheng et al. (2023) reaches the final answer by iteratively generating answers and using the previously generated answers as context hints.
@@
Self-consistency (Wang et al., 2023) refers to a simple decoding strategy for further improving reasoning performance, leveraging the fact that complex reasoning tasks typically allow for more than one correct reasoning path. Li et al. (2024) assign appropriate weights for answer aggregation to achieve adaptive self-consistency. Jain et al. (2023) and Wang et al. (2024) extend it for open-ended generation tasks like code generation and text summarization. However, they require multiple sampling with the pre-set size, which will incur much more computation cost. To realize cost-efficient self-consistency, Aggarwal et al. (2023) introduce an adaptive stopping criterion based on the amount of agreement between the samples so far. Li et al. (2023) divides the large preset sample size into several sequential small windows, and stop sampling when answers within a window are all the same.
@@ 5 Conclusion
In this work, we propose a novel cost-efficient decoding method called Difficulty-Adaptive Self-Consistency (DSC), which fully leverages the difficulty information of given problem set to allocate computational resources based on their respective levels of difficulty, so as to alleviate issues of redundant sampling on simple questions and multiple re-sampling limitations that current cost-efficient self-consistency methods face. Extensive experiments show that DSC consistently surpasses two strong cost-effcient self-consistency baselines ASC and ESC by a significant margin on six popular benchmarks, while attaining comparable accuracy. Additional experiments confirm the effectiveness of the proposed three-step algorithms: Difficulty Ranking, Problem Partition, and Sample Size Pre-Allocation.
@@
As illustrated in analysis of Difficulty Ranking, DSC is difficult to apply directly in scenarios that require real-time reasoning for mixed-type problems. Further classification of problems by type or further optimization of the current Difficulty Ranking algorithm is needed.

### v3

@@ Abstract
Self-consistency (SC), a widely used decoding strategy for chain-of-thought reasoning, shows significant gains across various multi-step reasoning tasks but comes with a high cost due to multiple sampling with the preset size. Its variants, Adaptive self-consistency (ASC) and Early-stopping self-consistency (ESC), dynamically adjust the number of samples based on the posterior distribution of a set of pre-samples, reducing the cost of SC with minimal impact on performance. Both methods, however, do not exploit the prior information about question difficulty. It often results in unnecessary repeated sampling for easy questions that could be accurately answered with just one attempt, wasting resources. To tackle this problem, we propose Difficulty-Adaptive Self-Consistency (DSC), which leverages the difficulty information of batch queries from both prior and posterior perspectives to adaptively allocate inference resources, further reducing the overall cost of SC. To demonstrate the effectiveness of DSC, we conduct extensive experiments on three popular categories of reasoning tasks: arithmetic, commonsense and symbolic reasoning on six benchmarks. The empirical results show that DSC consistently surpasses the strong baseline ASC and ESC in terms of costs by a significant margin, while attaining comparable performances.111Our code and data have been released on https://github.com/WangXinglin/DSC.
@@
Xinglin Wang1, Shaoxiong Feng2, Yiwei Li1, Peiwen Yuan1, Yueqi Zhang1, Chuyi Tan1, Boyuan Pan2, Yao Hu2, Kan Li1222Corresponding author. 1 School of Computer Science, Beijing Institute of Technology 2 Xiaohongshu Inc {wangxinglin,liyiwei,peiwenyuan, zhangyq, tanchuyi, likan}@bit.edu.cn {shaoxiongfeng2023}@gmail.com {panboyuan,xiahou}@xiaohongshu.com
@@ 1 Introduction
However, both approaches still suffer from the following two shortcomings: (1) Both require a certain amount of pre-sampling for all problems, which still results in redundant waste. As shown in Figure 1, there is a high overlap of inference results between SC and CoT, which suggests that only a small portion of questions (3.8% for GPT-4) benefit from SC. Generating multiple samples for the remaining questions compared to CoT (only sampling once) results in a significant waste of costs. Although ASC and ESC somewhat reduce the ineffective cost of SC by decreasing average sample size, there remains redundant sampling for simple problems222At least 3 samples per problem for ASC and 4 for ESC.. (2) Multiple re-samples bring additional significant input costs. Both ASC and ESC focus solely on reducing the number of outputs, without considering the extra input costs brought by multiple re-samples (Table 1). Consequently, for problems requiring multiple sampling times, ASC and ESC will introduce substantial extra input costs, sometimes outweighing the savings achieved through output sampling reduction (Table 3).
To alleviate these issues, we take inspiration from the strategies humans employ when solving reasoning problems within a limited total time. As illustrated in Figure 2, humans pre-assess the difficulty of the problems before reasoning, and adaptively allocating problem-solving time based on the accessed difficulty. In light of this, we propose Difficulty-Adaptive Self-Consistency (DSC), a novel method which leverages the difficulty information from both prior and posterior perspectives to further reduce the inference computational cost of SC. As illustrated in Figure 3, DSC consists of three steps. Firstly, we propose Difficulty Ranking algorithm which utilizes the LLM itself to rank the difficulty of the given problem set. While using LLM for ranking will incur additional costs, later experiments will demonstrate that the prior difficulty information gained from this process can lead to a greater reduction in overall inference costs. Based on this, we propose a Problem Partition strategy that divides the problem set into two parts, easy and hard, using the information regarding difficulty rankings. For problems belonging to easy part, a single CoT sampling is performed, which saves cost without sacrificing performance. Lastly, we propose Sample Size Pre-Allocation algorithm to predict the sample size needed for problems belonging to hard part, which reduce the number of re-samples and save the cost of inputs.
@@
We conducted extensive experiments on six popular benchmarks and validated the effectiveness and generalizability of our proposed method.
To the best of our knowledge, we are the first to propose dynamically allocating sampling resources based on query difficulty.
@@ 2.1 Difficulty Ranking
As problems usually do not carry difficulty label themselves, we seek to utilize the powerful comparative and ranking capabilities of LLM to obtain the difficulty rank of the given problem set. Considering the limited context window size of LLM and its lost in the middle (Liu et al., 2024b) issue in understanding long contexts, we randomly divide the problem set of into batches of size () and let LLM rank the difficulty of the problems in each batch. Since a single random split only allows us to obtain the difficulty ranking of each problem within the corresponding batch, we perform times random batch splitting, and sort the average difficulty rankings of each problem in different batches to obtain the difficulty ranking of entire problem set. Algorithm 1 illustrates the specific implementation of difficulty ranking. Specifically, is a dictionary that stores the difficulty levels of all questions. Its keys represent the indices of the questions, and its values are lists used to store the difficulties of the questions across different batches. Merge indicates appending the difficulty of question in the current batch to (List). Average refers to calculating the average value of , so as to obtain difficulty value of question i across the entire problem set. Given the additional costs of using LLM for ranking, we carefully design the ranking prompts333See Appendix A.1 for corresponding prompts. to minimize these costs, and later experiments will prove that the prior difficulty information gained can significantly reduce overall inference costs.
@@ 2.2 Problem Partition
After obtaining the problem set with difficulty ranking, we aim to find which part of the problems only require LLM to perform CoT sampling once. A simple and intuitive idea is that when LLM is very confident about a number of continuous problems sorted by difficulty (the results of all pre-samples are the same), it can be inferred that the problems easier than these problems only need one CoT sampling. Guided by this, we design the Problem Partition algorithm, illustrated in Algorithm 2. Specifically, we pre-sample the problem set sorted by difficulty from hard to easy, and store the entropy of pre-sampling results of each problem () in a list. We stop pre-sampling when the latest items in the list are all zero. The remaining problems without pre-sampling are then divided into the easy part, and a single CoT sampling is performed for each.
@@ 2.3 Sample Size Pre-Allocation
After performing CoT sampling on the easy part problems, we seek to predict and allocate the sample size for the problems belonging to hard part, so as to mitigate the substantial cost brought by multiple re-samples. Considering that the required sample size for each problem should be similar to those needed for problems of comparable difficulty, we predict the sample size of the current problem based on the total sample size of the nearest easier444As the pre-allocated sample size could be larger than the required sample size (causing higher output costs), we use the sample size of easier rather than harder neighboring problems for prediction. problems. Algorithm 3 shows the workflow of Sample Size Pre-Allocation. Specifically, we sample questions belonging to the hard part from easy to difficult. For the current question, we predict its pre-allocation sample size based on the average total sample size of its previous questions. Then, we judge the distribution of the current samples based on the stopping criteria 555Following ASC, we use Dirichlet Stopping Criteria as default. See Appendix A.8 for more details.. When the criteria is not met, we re-sample based on the expansion window , until the sampling distribution meets the criteria for stopping sampling or the number of samples reaches the max sample size . After sampling for the current question ends, we add its samples and total sample size to and lists respectively, in order to pre-allocate the sample size for the next question.
@@ 3.1.3 Implementation details
| Method | Input tokens | Output tokens | Cost | Acc |
| SC | 846 | 6045 | 0.3881 | 73.19 |
| ASC | 15982 | 3265 | 0.6754 | 73.14 |
| ESC | 4458 | 4415 | 0.3986 | 73.18 |
| DSC | 1617 | 3857 | 0.2799 | 73.18 |
@@ 3.2 Main Results
Table 2 summarizes the cost and accuracy of SC, ASC, ESC, and proposed DSC for each dataset. We show that DSC consistently outperforms all baselines on cost by a significant margin across all datasets, while barely affecting performance. Specifically, DSC reduces the cost on GPT-4 by an average of 65.29% and on GPT-3.5-Turbo by 56.04% compared to SC. In comparison to the strongest baseline method ESC, DSC reduces the cost on GPT-4 by an average of 24.81% and on GPT-3.5-Turbo by 21.86%, which demonstrates the effectiveness of DSC. Furthermore, the above results show that the prior difficulty information gained from difficulty ranking can help greatly lower overall inference costs, even though it first introduces some additional costs.
@@
DSC Maintains Generalizability on Small Open-source Language Model.
To further explore the effectiveness of DSC on small open-source models, we conduct experiments on the open-source model Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) using GSM8K dataset. As shown in Table 4, we convert the token cost into price according to that of GPT-4 for a simpler comparison and keep the other settings completely consistent with the main experiment777Please note that for DSC, the cost of all three sub-steps are included.. The experimental results indicate that DSC has the potential to work effectively on smaller models.
@@ 3.3.1 Difficulty Ranking
As we instruct LLM to produce extremely concise outputs for Difficulty Ranking, the cost of output tokens brought by Difficulty ranking is very low888With an average consumption of 15 tokens for each question.. Table 7 gives comparison of input tokens brought by Difficulty Ranking and few-shot reasoning across six datasets. We show that the average input cost for each problem’s difficulty ranking is less than the average single input cost for its few-shot reasoning. Considering that ESC and ASC require multiple re-samples for reasoning (multiple few-shot inputs), this additional input cost of Difficulty Ranking is totally acceptable.
@@ 3.3.2 Problem Partition
Regarding the intuitive idea proposed in Section 2.2, a straightforward question is how large the judge window size needs to be to ensure the accuracy on simpler problems with single CoT sampling. To answer this, we conduct an experiment of DSC under different judge window size with GPT-4 on GSM8K. As shown in Figure 4, the accuracy increases with the rise of until it reaches 32. Meanwhile, the cost of DSC also increases with the enlargement of , as the easy part divided by the problem partition decreases with its increase. Therefore, it is a good choice to set a smaller judge window size while ensuring accuracy.999We select 32 as the default value for .
@@ 3.3.3 Sample Size Pre-Allocation
Given that Sample Size Pre-Allocation can trade output tokens for savings in input tokens, a natural question is when the max total savings can be achieved. As shown in Figure 5, we count the tokens of input and output and the corresponding cost of DSC under different prediction window size . The experimental results show that the cost of DSC decreases rapidly first and then slighly increases with the increase of , and the input tokens also follow a similar trend. As rises, simpler questions predict larger sample sizes for the current query. However, when becomes sufficiently large, the predicted sample size falls below the actual value, leading to more re-sampling. In contrast, output tokens gradually decrease because, as increases, the predicted sample size continues to decline, reducing redundancy in output token sampling. In summary, it is suggested to select within a relatively moderate range (e.g. from 8 to 32).
@@ 4 Related Work
Chain-of-thought prompting has been proven to be an effective method of solving complex reasoning problems (Wei et al., 2022a). By following the pattern of gradually solving sub-problems, few-shot CoT (Fu et al., 2023) are capable of stimulating LLM reasoning abilities. On this basis, Least-to-most prompting (Zhou et al., 2023) suggest explicitly splitting the problem and solving them step by step. Zheng et al. (2023) reach the final answer by iteratively generating answers and using the previously generated answers as context hints.
@@
Self-consistency (Wang et al., 2023) refers to a simple decoding strategy for further improving reasoning performance, leveraging the fact that complex reasoning tasks typically allow for more than one correct reasoning path. Li et al. (2024) assign appropriate weights for answer aggregation to achieve adaptive self-consistency. Jain et al. (2023) and Wang et al. (2024) extend it for open-ended generation tasks like code generation and text summarization. However, they require multiple sampling with the pre-set size, which will incur much more computation cost. To achieve cost-efficient self-consistency, Aggarwal et al. (2023) introduce an adaptive stopping criterion based on the amount of agreement between the samples so far. Li et al. (2023) divide the large preset sample size into several sequential small windows, and stop sampling when answers within a window are all the same.
Difficulty-adaptive Test-Time Scaling
Test-time scaling (Wu et al., 2024a; Snell et al., 2024; Brown et al., 2024) has been widely explored as an effective approach to improving model performance. However, conventional test-time scaling applies a uniform amount of computation to all queries, regardless of their difficulty, resulting in unnecessary computational overhead. To mitigate this inefficiency, recent work has proposed difficulty-adaptive scaling strategies that dynamically allocate computational resources based on query difficulty. Snell et al. (2024) validate the feasibility of optimizing test-time scaling compute based on query difficulty. To address excessive refinement in self-refinement, Chen et al. (2024) classify problems as easy or hard, solving easy ones with coarse aggregation and hard ones with fine-grained, iterative multi-agent refinement. Manvi et al. (2024) introduce capability-aware and mid-generation self-evaluations to enable difficulty-aware adaptive test-time computation by dynamically allocating compute based on query difficulty and model capability. Similarly, Damani et al. (2024) train a lightweight model to estimate query difficulty and allocate extra computation to those queries that would benefit most from a more computationally intensive decoding procedure.
@@ 5 Conclusion
In this work, we propose a novel cost-efficient decoding method called Difficulty-Adaptive Self-Consistency (DSC), which fully leverages the difficulty information of given problem set to allocate computational resources based on their respective levels of difficulty, so as to alleviate issues of redundant sampling on simple questions and multiple re-sampling limitations that current cost-efficient self-consistency methods face. Extensive experiments show that DSC consistently surpasses two strong cost-effcient self-consistency baselines ASC and ESC by a significant margin on six popular benchmarks, while attaining comparable accuracy. Additional experiments confirm the effectiveness of the proposed three-step algorithms.
@@
As the performance of difficulty ranking declines when handling mixed-type problems (Table 10), DSC may encounter challenges when applied in scenarios requiring real-time reasoning for mixed-type problems. Exploring further classification of problems by type or optimizing the current Difficulty Ranking algorithm could enhance its applicability.
@@
Ethics Statement
All of the datasets used in this study were publicly available, and no annotators were employed for our data collection. We confirm that the datasets we used did not contain any harmful content and was consistent with their intended use (research). We have cited the datasets and relevant works used in this study.
Acknowledgements
This work is supported by the Beijing Natural Science Foundation, China (Nos. 4222037, L181010).